OpenStudy (anonymous):

Devon and his friends bought strawberry wafers for $3 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $30 to buy a total of 22 packets of wafers of the two varieties. Write a system of equations that can be solved to find the number of packets of strawberry wafers and the number of packets of chocolate wafers that Devon and his friends bought at the carnival. Define the variables used in the equations.

OpenStudy (kash_thesmartguy):

What do you think?

OpenStudy (anonymous):

i dont know

OpenStudy (kash_thesmartguy):

here, we have to make two equations. So, for the first equation, if they bought x number of strawberry wafers and y number of chocolate wafers, how much wafers did they buy (Remember -> just an equation.)

OpenStudy (kash_thesmartguy):

The first equation will look something like this: \[s+c=22\]

OpenStudy (kash_thesmartguy):

You getting this, right? @hayespurdy

OpenStudy (anonymous):

yes

OpenStudy (kash_thesmartguy):

For number 2 equation, it should be how much they spent. So: \[3x+1y=$30\]And there all of the answers! The two equations are your answer

OpenStudy (kash_thesmartguy):

BTW, can I get a medal? @hayespurdy

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

What about How many packets of chocolate wafers and strawberry wafers did they buy? Explain how you got the answer and why you selected a particular method to get the answer @Kash_TheSmartGuy

OpenStudy (kash_thesmartguy):

3x+1y=30, solve that and there's your answer for how many wafers. I chose this method because it was easy, productive and fast.

OpenStudy (anonymous):

how do u solve it @Kash_TheSmartGuy